# Jaisidh Singh1,2,4, Diganta Misra3,4, Boris Knyazev5, Antonio Orvieto3, 4, 6

### v1

@@ Abstract
Foundation multi-modal models are often designed by stitching of multiple existing pretrained uni-modal models: for example, an image classifier with an autoregressive text model. This stitching process is performed by training a connector module that aims to align the representation-representation or representation-input spaces of these uni-modal models. However, given the complexity of training such connectors on large scale web-based datasets coupled with the ever-increasing number of available pretrained uni-modal models, the task of uni-modal models selection and subsequent connector module training becomes computationally demanding. To address this under-studied critical problem, we propose Hypernetwork Model Alignment (Hyma), a novel all-in-one solution for optimal uni-modal model selection and connector training by leveraging hypernetworks. Specifically, our framework utilizes the parameter prediction capability of a hypernetwork to obtain jointly trained connector modules for $N\times M$ combinations of uni-modal models. In our experiments, Hyma reduces the optimal uni-modal model pair search cost by 10×10$\times$10 × (averaged across all experiments), while matching the ranking and trained connector performance obtained via grid search across a suite of diverse multi-modal benchmarks.
@@ 1 Introduction
  • Pairing: Given a set of $N$ models in modality 1 (e.g., vision) and $M$ models in modality 2 (e.g., text), provide the optimal (best performing) combination pair ($n,m;\ n\in N\ |\ m\in M$) to construct a multi-modal model for a target task and/or under target constraints (e.g., parametric size, embedding dimensions, latency).
@@
We present Hypernetwork Model Alignment (Hyma), a method that, given $N$ modality 1 (e.g., image) and $M$ modality 2 (e.g., text) models, leverages a hypernetwork Ha et al. (2016) that jointly learns to generate connectors for all possible $N\times M$ combinations. Our approach serves both as an indicator for the optimal model pair configuration and also as a trainer that produces stitched multi-modal models that perform on par with the best stitched model pair obtained via exhaustive grid search. In our experiments, where $N\times M$ can be as high as 27 (as discussed in Section 5), this method enables an efficiency gain of 10×10$\times$10 × in obtaining the best stitched model pair compared to grid search.
@@ 4.1 Baselines
  • AutoPair: To enable a fair comparison with Hyma’s efficiency-focused design, we implement an pairing baseline that iteratively searches a given set of pairs by training for a fixed number of epochs, and then prunes all pairs below the median performance. Specifically, AutoPair optimizes model pair selection and stitching within a FLOPs budget equal to that used by Hyma for the same model zoo. Full implementation details are provided in Section 5.4.
@@ 5 Empirical Results
For multi-modal image classification on the ImageNet-1K, we find that the ranking order of the stitching performed by Hyma reflects that found by full grid search to strong extent. This is indicated by the normalized discounted cumulative gain (NDCG @$\ k$) computed for the top 5 and 7 ranks. Additionally, Spearman’s $\rho$ across all $N\times M=27$ ranks further corroborates this. Notably, both NDCG @$\ k$ and Spearman’s $\rho$ for CIFAR-100 are lower in value w.r.t ImageNet-1K. In terms of performance gains, Hyma improves upon random selection of encoder pairs to stitch, as well as selecting encoders based on their uni-modal performance. Interestingly, we find that asking a massively pretrained LLM such as Claude 4 Sonnet yields a similar result to UniT-1. For Oracle (GS), find that the best stitchings generated by Hyma underperform average of $4.84\ \%$ and $2.88$ for ImageNet-1K and CIFAR-100 across all connector types. However, this occurs at $\times1$ fewer FLOPs spent.
@@
Table 4: Hyma vs AutoPair Results: We show the performance gain ($\Delta$) of the best connector (for all connector configurations) obtained post stitching via Hyma, compared to that obtained via AutoPair. This is done for a model zoo of $N\times M=12$.
@@ 5.3 Connection to Data Pruning
Method Best Model configuration Perf.
C-GS DeiT-3S + miniLM-L 24.07
HYMA DeiT-3S + miniLM-L 27.46
Table 5: HYMA vs. Constrained Grid Search (C-GS). For the setting $N\times M=3$, $B_{m}=1$, we constrain the total data available to Grid Search to one-third, aligning it with HYMA’s data budget. While this constraint results in a comparable reduction in FLOPs relative to full Grid Search, it leads to a notable drop in performance. Perf. denotes MIC top-1 accuracy on the ImageNet-1k dataset.
5.4 AutoPair
We conduct a step-wise search-and-prune procedure over $6$ image encoders (evenly split across embedding dimensions $768$ and $1024$) and $2$ text encoders (also evenly split across embedding dimensions $768$ and $1024$). First we initialize a FLOPs budget equal to the total FLOP cost of searching over $N\times M=27$ pairs with Hyma for $10$ epochs. Next, our procedure trains connectors between all $12$ pairs for 2 epochs each, after which we rank each connector by its performance on a given task and dataset. After the ranking, we prune all pairs that exhibit performance that is less than or equal to the median performance. This is repeated until we exhaust the budget. If we are left with only one model after iterative pruning, we train it until the budget is exhausted.
@@ Acknowledgements
Table 6: Hyma MLLM Results: We report the ranking similarity between Hyma and the Oracle—Grid Search (GS)—using NDCG and Spearman’s $\rho$. Across all three connector configurations, Hyma exhibits strong correlation with GS rankings. Additionally, we show the perplexity difference ($\Delta$) of the best connector obtained post stitching via Hyma, compared to four baselines: (a) Random, random pairing and stitching (avg. over 5 runs); (b) UniT-1, stitching the best unimodal models; (c) Ask-LLM, model pairs picked by GPT-4.1; and (d) Oracle, grid search over all $N{\times}M{=}9$ configurations.
@@ B.1 Image encoders (source: timm Wightman (2019))
Table 8: All pretrained text encoders and LLMs used in our work are given above, along with their shorthand IDs that may be referred to in the main manuscript and the appendix.
@@ Appendix C Designing the Model Zoo
While our empirical analysis suggests that models with larger parametric capacity or higher embedding dimensionality generally perform better after stitching, a natural question arises: why include smaller models in the model zoo at all? We justify their inclusion based on the following two considerations:
@@
    Second, our empirical observations indicate that larger models are not always the best-performing choice when stitched into multi-modal pairs. This motivates the inclusion of a diverse set of model configurations in our zoo to better explore the multi-modal design space. By covering a broader range of capacity combinations, HYMA facilitates a more comprehensive and efficient search. This point is further supported by the observations made in Figure 1 and Table 1.
@@ Appendix D Training and hyper-parameter details
For MLLMs, we follow recipes given in Jia et al. (2024) for training only the connector (referred to as the feature alignment phase of pretraining). Particularly, we use a batch size of $64$ for training individual connectors and a learning rate of $31e-$ 3. This is subject to a schedule of warmup ratio $23e-$ 2 following a cosine decay to 0. The batch size training Hyma for MLLMs is $32$ and the learning rate is $31e-$ 3. We choose our optimizer as Adam.
@@ G.1 Ask-LLM (for VLMs)
We prompt Claude 4 Sonnet to identify the best model pair:

### v2

@@ Abstract
Foundation multi-modal models are often designed by stitching of multiple existing pretrained uni-modal models: for example, an image classifier with an text model. This stitching process is performed by training a connector module that aims to align the representation spaces of these uni-modal models towards a multi-modal objective. However, given the complexity of training such connectors on large scale web-based datasets coupled with the ever-increasing number of available pretrained uni-modal models, the task of uni-modal models selection and subsequent connector module training becomes computationally demanding. To address this under-studied critical problem, we propose Hypernetwork Model Alignment (Hyma), a novel all-in-one solution for optimal uni-modal model selection and connector training by leveraging hypernetworks. Specifically, our framework utilizes the parameter prediction capability of a hypernetwork to obtain jointly trained connector modules for $N\times M$ combinations of uni-modal models. In our experiments, Hyma reduces the cost of searching for the best performing uni-modal model pair by 10×10$\times$10 ×, while matching the ranking and trained connector performance obtained via grid search across a suite of diverse multi-modal benchmarks.
@@ 1 Introduction
  • Pairing: Given a set of $N$ models in modality 1 (e.g., vision) and $M$ models in modality 2 (e.g., text), provide the optimal (best performing) combination pair ($n,m;\ n\in N\ |\ m\in M$) to construct a multi-modal model for a target task and/or under target constraints (e.g., parametric size, embedding dimensions).
@@
We present Hypernetwork Model Alignment (Hyma), a method that, given $N$ modality 1 (e.g., image) and $M$ modality 2 (e.g., text) models, leverages a hypernetwork Ha et al. (2016) that jointly learns to generate connectors for all possible $N\times M$ combinations. Our approach serves both as an indicator for optimal model pair configurations and as a trainer that produces stitched multi-modal models performing on par with the best stitched model pair obtained via grid search. In our experiments, where $N\times M$ can be as high as 27 (discussed in Section 5), our method enables an efficiency gain of 10×10$\times$10 × in obtaining the best stitched model pair compared to grid search.
@@ 4.1 Baselines
  • AutoPair: To enable a fair comparison with Hyma’s efficiency-focused design, we implement an pairing baseline that iteratively searches a given set of pairs by training for a fixed number of epochs, and then prunes all pairs below the median performance. Specifically, AutoPair optimizes model pair selection and stitching within a FLOPs budget equal to that used by Hyma for the same model zoo. More details are provided in Section 5.3.
@@ 5 Empirical Results
For multi-modal image classification on the ImageNet-1K, we find that the ranking order of the stitching performed by Hyma reflects that found by full grid search to strong extent. This is indicated by the normalized discounted cumulative gain (NDCG @$\ k$) computed for the top 5 and 7 ranks. Additionally, Spearman’s $\rho$ across all $N\times M=27$ ranks further corroborates this. Notably, both NDCG @$\ k$ and Spearman’s $\rho$ for CIFAR-100 are lower in value w.r.t ImageNet-1K. In terms of performance gains, Hyma improves upon random selection of encoder pairs to stitch, as well as selecting encoders based on their uni-modal performance. Interestingly, we find that asking a massively pretrained LLM such as Claude 4 Sonnet yields a similar result to UniT-1. For Oracle (GS), find that the best stitchings generated by Hyma underperform average of $4.84\ \%$ and $2.88$ for ImageNet-1K and CIFAR-100 across all connector types. However, this occurs at 10×10$\times$10 × fewer FLOPs spent.
@@
Table 4: Hyma vs AutoPair Results ($N\times M=12$): We show the performance gain ($\Delta$) of the best connector (for all connector configurations) obtained post stitching via Hyma, compared to that obtained via AutoPair.
@@
5.3 Hyma vs AutoPair
We conduct a step-wise search-and-prune procedure over $6$ image encoders (evenly split across embedding dimensions $768$ and $1024$) and $2$ text encoders (also evenly split across embedding dimensions $768$ and $1024$). First we initialize a FLOPs budget equal to the total FLOP cost of searching over $N\times M=12$ pairs with Hyma for $10$ epochs. Next, our procedure trains connectors between all $12$ pairs for 2 epochs each, after which we rank each connector by its performance on a given task and dataset. After the ranking, we prune all pairs that exhibit performance that is less than or equal to the median performance. This is repeated until we exhaust the budget. If we are left with only one model after iterative pruning, we train it until the budget is exhausted.
As shown in Table 4, stitches obtained by AutoPair exhibit significantly lower performance than those obtained via Hyma, as the budget finishes before the individually trained connectors can reach strong performance.
@@ 5.4 Connection to Data Pruning
Method Best Model configuration Perf.
C-GS DeiT-3S + miniLM-L 24.07
HYMA DeiT-3S + miniLM-L 27.46
Table 5: Hyma vs. Constrained Grid Search (C-GS). For the setting $N\times M=3$, $B_{m}=1$, we constrain the total data available to Grid Search to one-third, aligning it with HYMA’s data budget. While this constraint results in a comparable reduction in FLOPs relative to full Grid Search, it leads to a notable drop in performance. Perf. denotes MIC top-1 accuracy on ImageNet-1K.
@@ Acknowledgements
Table 6: Hyma MLLM Results: We report the ranking similarity between Hyma and the Oracle—Grid Search (GS)—using NDCG and Spearman’s $\rho$. Across all three connector configurations, Hyma exhibits strong correlation with GS rankings. Additionally, we show the perplexity difference ($\Delta$) of the best connector obtained post stitching via Hyma, compared to four baselines: (a) Random: random pairing and stitching (avg. over 5 runs); (b) UniT-1: stitching the best unimodal models; (c) Ask-LLM: model pairs picked by Claude 4 Sonnet; and (d) Oracle: Full grid search over all $N{\times}M{=}9$ configurations.
@@ B.1 Image encoders (source: timm Wightman (2019))
Table 8: All pretrained text encoders and LLMs used in our work are given above, along with their shorthand IDs that may be referred to in the main manuscript.
@@ Appendix C Designing the Model Zoo
While our empirical analysis suggests that models with larger parametric capacity or higher embedding dimensionality generally perform better after stitching, a natural question arises: why include smaller models in the model zoo at all? We justify their inclusion based on the following:
@@
    Second, our empirical observations indicate that larger models are not always the best-performing choice when stitched into multi-modal pairs. This motivates the inclusion of a diverse set of model configurations in our zoo to better explore the multi-modal design space. By covering a broader range of capacity combinations, HYMA facilitates a more comprehensive and efficient search, supported by observations from Figure 1 and Table 1.
@@ Appendix D Training and hyper-parameter details
For MLLMs, we follow recipes given in Jia et al. (2024) for training only the connector (referred to as the feature alignment phase of pretraining). Particularly, we use Adam with batch size of $64$ for training individual connectors and learning rate $31e-$ 3. This is subject to a schedule of warmup ratio $23e-$ 2 following a cosine decay to 0. The batch size training Hyma for MLLMs is $32$ and the learning rate is $31e-$ 3.
@@ G.1 Ask-LLM (for VLMs)
We prompt Claude 4 Sonnet Anthropic to identify the best model pair: